### v1

# Decoherence in Exchange-Coupled Quantum Spin Qubit Systems: Impact of Multiqubit Interactions and Geometric Connectivity

Abstract
We investigate the impact of different connectivities on the decoherence time in quantum systems under quasi-static Heisenberg noise. We considered three types of fundamental units, including node, stick and triangle and connect them into rings, chains, and trees. We find that rings exhibit greater stability compared to chains, contrary to the expectation that higher average connectivity leads to decreased stability. Additionally, the “stick” configuration is more stable than the “triangle” configuration. We also observe similar trends in entanglement entropy and return probability, indicating their potential use in characterizing decoherence time. Our findings provide insights into the interplay between connectivity and stability in quantum systems, with implications for the design of robust quantum technologies and quantum error correction strategies.
I Introduction
To pursue a practical quantum computer, it is necessary to successfully combine numerous qubits and ensure high levels of accuracy in single and two-qubit manipulations, as quantum error correction requires $>99.9$% fidelity in gate operations. Among different qubit platforms, while ultracold atoms and Si spin qubits are most promising ones for achieving scalability, operations on these systems suffer from lower fidelity rates as compared to ion traps and superconducting qubits Ladd et al. (2010); Zhang et al. (2017); Bloch (2008); Schneider and Saenz (2012); Levine et al. (2019); Monroe and Kim (2013); Keith et al. (2022). While previous experimental studies have demonstrated high fidelity in operations on single Yang et al. (2019); Gilbert et al. (2020); Chan et al. (2021) or two-qubit systems Veldhorst et al. (2014); Huang et al. (2019); Zajac et al. (2018); Watson et al. (2018); Xue et al. (2019); Mills et al. (2022); Noiri et al. (2022); Xue et al. (2022), achieving the same level of fidelity in multiqubit systems remains challenging as the scale of the system increases. Improving scalability is therefore a key problem in the study of the physical realization of quantum information processing, relevant to various platforms including dopants in semiconductors, gate-defined quantum dots, photons and atoms in cavities, Rydberg atom arrays, superconducting quantum circuits, and trapped atomic ions Britton et al. (2012); Bohnet et al. (2016); Altman et al. (2021).
To improve the fidelity of multiqubit systems, it is important to investigate the factors that may impact their decoherence, and subsequently, the fidelity of applied quantum controls as the system size increases Reagor et al. (2018); Fang et al. (2023); Kim et al. (2022). In order to execute quantum algorithms, qubits are necessarily connected together with possible operations on the connected ones. However, such connectivity usually implies decoherence, as can be seen in the case of ultracold atom systems and IBM’s superconducting qubits Bloch (2008); Zhang et al. (2017). In a quantum device, two aspects on how the qubits are connected worth consideration: the number of links between the pairs of qubits, as well as how the qubits are geometrically connected. Intuitively, if there are less connections between qubits, the device would have longer coherence time because the qubits are less affected by their neighbors. However, as shown in an important prior work Throckmorton and Das Sarma (2022), such perception is not entirely accurate and the geometry of the device, such as ring and chain configuration, is also important. It has been shown that larger average connectivity does not necessarily lead to faster decay of coherence. Nevertheless, the effect of geometric connectivity on decoherence time remains an open question. Therefore, in this work, our aim is to investigate more comprehensively into this aspect and provide a more thorough understanding on how the geometric connectivity affects decoherence in multiqubit systems.
In this paper, we investigate the impact of different geometric connectivity patterns using various elementary units on the decoherence time under quasi-static Heisenberg noise. We begin by introducing a range of elementary units and then assemble them into ring, chain, and tree configurations to determine which arrangement can maintain quantum state stability for a longer duration as the system size increases. At the first glance, one may have the perception that more links between qubits leads to less stability, provided that the system size is fixed. Surprisingly, we find that the geometric connectivity plays a key role, such that the ring configuration is generally more stable than the chain, even if with a greater number of links. Furthermore, we analyze the entanglement entropy and observe similar trends to the return probability in both chain and tree configurations. These similarities suggest that both entanglement entropy and return probability can be effectively utilized to characterize the decoherence time.
The rest of the paper is organized as follows. In Sec. II, we introduce our model and describe the elementary units and geometric connectivity configurations, including rings, chains, and trees. We also outline the method used to extract the decoherence time. Moving on to Sec. III, we present the results of our study, comparing the decoherence time of elementary units (“node”, “stick”, and “triangle”) in ring and chain configurations (Sec. III.1). Additionally, we analyze the entanglement entropy and its relationship to the return probability for ring, chain, and tree configurations (Sec. III.2). Finally, in the concluding section, Sec. IV, we summarize our findings and their implications, highlighting the importance of geometric connectivity in understanding and mitigating decoherence effects in quantum systems.
II Model and method
In this study, we investigate the impact of geometric connectivity on decoherence time in a system with Heisenberg interaction. We consider different connectivity patterns, including ring, chain, and tree configurations constructed from multiple elementary units.
We start with a system with Heisenberg interaction on connected qubits, described by the Hamiltonian:
$H=\sum_{ij}J_{ij}(t)\bm{\sigma}_{i}\cdot\bm{\sigma}_{j},$ (1)
where $\bm{\sigma}_{i}$ is the vector of Pauli matrices on spin $i$. Experimental evidence suggests that the Heisenberg interaction is the dominant source of noise between qubits, denoted by $J_{ij}(t)$. By studying the evolution of decoherence time under different connectivity configurations, we aim to gain a comprehensive understanding of the relationship between geometric connectivity and decoherence in multiqubit systems.
In the presence of Heisenberg interaction, it is important to consider both static and dynamic aspects of the noise. In experimental settings, the frequency of noise fluctuations is much lower than the characteristic timescale of the experimental manipulations. Therefore, we primarily focus on the effect of quasi-static noise, which is characterized by a normal distribution. To account for this noise, we employ the Monte Carlo method to randomly select noise intensities and average their effects. Specifically, for quasi-static noise, we model the coupling strength $J_{ij}(t)$ as a Gaussian distribution with a mean value $J_{ij,0}$ and a deviation $\sigma$. It is worth noting that as experimental operations become more accurate in the future, the influence of dynamic noise may become more significant. In this paper, our main focus is on the study of quasi-static noise and its impact. The effects of dynamic noise are discussed in Appendix A.
Next, we incorporate the study of geometric connectivity into our model. We consider several different elementary units, including the “node”, “stick”, and “triangle”, and explore the geometric connectivity configurations of ring, chain, and tree. As depicted in Fig. 1, the vertices represent qubits, the gray lines represent the parts within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. These elementary units are then geometrically connected in different configurations, in the forms of a ring, chain, or tree, as shown in Fig. 2. In this study, we investigate the decoherence time of these elementary units with $L$ ranging between 4 and 10, where $L$ represents the number of qubits.
Figure 1: The figure illustrates the three different elementary units used in our study: “node”, “stick”, and “triangle”. The vertices represent qubits, the gray lines indicate the regions within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. Our focus in this paper is to investigate the effects of different geometric connectivity configurations on the system.
Figure 2: Illustration of chain, ring and tree configurations using different elementary units. The vertices represent qubits, the gray lines indicate the regions within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. The arrows inside the vertices represent the initial states that we choose. An upward arrow indicates a spin-up state at this site, while a downward arrow indicates a spin-down state. In our study, we focus on the qubits that are affected by noise-induced decoherence, which are the qubits between which quantum gates are applied. Moreover, only connected qubits can implement two-qubit gates. Therefore, we adopt a model with Heisenberg interaction between connected qubits to investigate the impact of noise.
We extract the decoherence time by calculating the return probability
$P(t)=\left|\langle\Psi(0)|\Psi(t)\rangle\right|^{2},$ (2)
where $|\Psi(0)\rangle$ is the initial state and $|\Psi(t)\rangle$ is the state at time $t$. We choose the initial state $|\Psi(0)\rangle=|\uparrow\downarrow\cdots\rangle$, as shown in Fig. 2. The return probability $P(t)\propto P_{A}(t)P_{\phi}(t)$ can be expressed as
$P_{A}(t)=P_{\infty}\pm(1-P_{\infty})e^{-(t/T_{2}^{*})^{\alpha}},$ (3)
where $P_{A}(t)$ is the amplitude term, $P_{\phi}(t)$ is the frequency term, and the decoherence time $T^{*}_{2}$ is extracted from fitting. In the case of elementary unit “node” with quasi-static noise, we analyze the probability $P(t)$ as a function of time $t$. Fig. 3 shows the evolution of $P(t)$ for $L=6,J_{0}=100/t_{0},\sigma=0.4/t_{0}$, represented by the solid line. In this work, we denote $t_{0}$ as our time unit, which is about 10 ns in superconducting qubits and about 1 $\mu s$ for a typical iron trap quantum device. The dashed line corresponds to the fitted curve, while the dotted line represents the asymptotic value $P_{\infty}$. We extract the envelope lifetime using the Eq. (3). The obtained decoherence time is $T_{2}^{*}/t_{0}=0.489$. This quantity is commonly measured in experimental settings and provides valuable insights into the system’s decoherence dynamics.
Figure 3: The probability $P(t)$ as a function of time $t$ for elementary unit “node”, with quasi-static noise under chain configuration. The solid line represents the evolution of $P(t)$ as a function of time under $L=4,J_{0}=100/t_{0},\sigma=0.4/t_{0}$, $t_{0}$ is the time unit, while the dashed line corresponds to the fitted curve. We extract the envelope lifetime based on Eq. (3) to characterize the decoherence time, and we obtain $T^{*}_{2}=0.706/t_{0}$.
III Results
III.1 Decoherence time under quasi-static noise
We initially focus on the ring and chain connectivity configurations. We will discuss the tree configuration in a later stage because the system size involved may be significantly larger. We specifically investigate the results of the “node” elementary unit’s tree configuration for a system size of $L=8$ in Sec. III.2.
Firstly, we extract the decoherence time $T_{2}^{*}$ for the elementary unit “node”, with the system size up to $L=10$. We set the interaction strength as $J_{0}=100/t_{0}$ and the deviation as $\sigma=0.5/t_{0}$. Choosing $J_{0}=100/t_{0}$ allows for faster decay of the system and more accurate calculation of $T_{2}^{*}$, as the product of the energy unit $J_{0}$ and the decoherence time remains constant. Fig. 4 shows that the decoherence time $T_{2}^{*}$ exhibits similar behavior to that reported in Ref. Throckmorton and Das Sarma (2022). It can be observed that in the ring configuration, with the “node” as the elementary unit, the overall decoherence time decreases as the system size $L$ increases. Specifically, the decoherence time decreases from $T_{2}^{*}=1.004/t_{0}$ at $L=4$ to $T_{2}^{*}=0.201/t_{0}$ at $L=10$, resulting in a decline of approximately $80\%$. Moreover, the even-qubit systems exhibit longer lifetimes compared to the odd-qubit systems. On the other hand, for the chain configuration, the decoherence time also decreases with increasing system size $L$, but at a slower rate compared to the ring configuration. Similar results have been reported in Ref. Throckmorton and Das Sarma (2022). Specifically, for the chain, the decoherence time decreases from $T_{2}^{*}=0.543/t_{0}$ at $L=4$ to $T_{2}^{*}=0.297/t_{0}$ at $L=10$, corresponding to a decline of approximately $45\%$. Minor discrepancies between our numerical results and those of Ref. Throckmorton and Das Sarma (2022) may arise from differences in the chosen energy scale, variations in the Monte Carlo process, and numerical fluctuations when extracting $T_{2}^{*}$ from the envelope fitting.
Figure 4: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary unit “node”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decay trend is similar to Ref. Throckmorton and Das Sarma (2022).
Next, we investigate the behavior of elementary units “stick” and “triangle” up to $L=10$, considering an interacting strength of $J_{0}=100/t_{0}$ and a deviation of $\sigma=0.5/t_{0}$. In Fig. 5, we observe that for the ring geometry with the “stick” as the elementary unit, the overall decoherence time decreases as the system size $L$ increases. Specifically, the decoherence time decreases by $48\%$, from $T_{2}^{*}=1.185/t_{0}$ at $L=6$ to $T_{2}^{*}=0.614/t_{0}$ at $L=10$. We note that we excluded the case of $L=4$ for the “stick” elementary unit, as it comprises only two elementary units, which cannot form a ring. As for the chain geometry with the “stick” as the elementary unit, the overall decoherence time also decreases with increasing system size $L$. Comparatively, at $L=6$, the decoherence time decreases from $T_{2}^{*}=0.625/t_{0}$ to $T_{2}^{*}=0.444/t_{0}$ at $L=10$, representing a decrease of $29\%$. In summary, the chain exhibits a slower rate of decrease in decoherence time with increasing size. However, the ring geometry consistently displays greater stability than the chain at the same system size.
Figure 5: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary units “stick”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decoherence time for ring configuration is longer than that of chain configuration.
In Fig. 6, we observe consistent trends in the decoherence time for both ring and chain geometries. For the ring geometry with the “triangle” as the elementary unit, the decoherence time decreases gradually as the system size $L$ increases. Specifically, $T_{2}^{*}$ decreases from $0.655/t_{0}$ for $L=4$ to $0.414/t_{0}$ for $L=10$. On the other hand, in the chain geometry with the “triangle” as the elementary unit, the overall decoherence time also decreases with increasing system size $L$, showing a similar rate of decrease. The decoherence time decreases from $T_{2}^{*}=0.462/t_{0}$ for $L=4$ with to $T_{2}^{*}=0.107/t_{0}$ for $L=10$. The overall rate of decay is relatively similar between the ring and chain geometries, highlighting the distinct influence of the elementary unit “triangle” and “stick” on the decoherence dynamics.
Figure 6: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary units “triangle”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decoherence time for ring configuration is also longer than that of chain configuration.
The average number of connections between qubits in a ring configuration consistently exceeds that found in a chain configuration. This observation leads to an intuitive hypothesis: a lower average number of qubit connections could potentially extend the decoherence time. The rationale behind this is straightforward — fewer qubit connections imply limited positions for the application of quantum gates, which in turn may reduce the impact of noise effects on the system. This relationship between the complexity of qubit interconnections and decoherence dynamics merits further exploration, as it could offer insights into optimizing quantum computing architectures for enhanced stability. However, in Fig. 5 and Fig. 6, we can conclude that the configuration ring is in general more stable than chain.
Now, we proceed to calculate the decoherence time $T_{2}^{*}$ as a function of the system size $L$ for the elementary units “stick” and “triangle” with ring and chain configurations, considering different deviation values $\sigma$.
(a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption
Figure 7: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for (a) elementary units “stick” with ring configuration , for (b) elementary units “stick” with chain configuration, for (c) elementary units “triangle” with ring configuration, and for (d) elementary units “triangle” with chain configuration with different deviation $\sigma$ and interacting strength $J_{0}=100/t_{0}$.
In Fig. 7(a) and Fig. 7(b), we present the decoherence times for the elementary units “stick” with ring and chain configuration, considering different deviation $\sigma$ and interacting strength $J_{0}=100/t_{0}$. Due to the limitation of the elementary unit “stick”, the case of $L=4$ is excluded as it consists of only two elementary units, which cannot form a ring. Therefore, we focus on the $L=6$ case for the ring. Notably, in the ring configuration, the overall decoherence times are longer compared to the chain. The dependence of the ring and chain to system size varies with different $\sigma$. For relatively large $\sigma$ values, such as $\sigma=0.5/t_{0}$, the decoherence time for the ring configuration decreases faster as the system grows. However, for relatively small $\sigma$ values, e.g. $\sigma=0.2/t_{0}$, the decoherence time decreases slower as the system size increases. Nevertheless, regardless of the value of $\sigma$, the stability of the ring is consistently better to that of the chain at the same system size.
In Figures 7(c) and 7(d), we investigate the decoherence times of the elementary unit “triangle” for both ring and chain configurations, taking into account varying values of deviation $\sigma$ and interaction strength $J_{0}=100/t_{0}$. Similar to the “stick” configuration, the ring geometry consistently shows longer overall decoherence times compared to the chain. Unlike the “stick” configuration, however, the ring demonstrates weaker dependence to system size for the elementary unit “triangle”. This implies that, irrespective of the $\sigma$ value, the decoherence time in the ring configuration decreases at a slower rate as the system size increases. Nonetheless, as observed in earlier comparisons, the ring configuration continues to outperform the chain in terms of stability, maintaining longer decoherence times than the chain at the same system size.
Another observation from Fig. 7 is that the “stick” elementary unit consistently outperforms the “triangle” in terms of stability. For example, by comparing Fig. 7(a) and Fig. 7(c), or Fig. 7(b) and Fig. 7(d), we can clearly see that, under the same $\sigma$ and $L$ with the same geometric configuration, the decoherence time of the quantum state with the “stick” elementary unit is always longer than that with the “triangle” elementary unit. We conclude that the “stick” elementary unit displays greater stability compared to the “triangle” unit.
Regarding the observation that rings are more stable than chains, we speculate that this behavior may be attributed to symmetry. For instance, in a three-qubit system without considering variance, the Hamiltonian of the ring commutes with the total spin Buterakos and Das Sarma (2021), thereby conserving the total spin and providing some protection to the quantum state against evolving into other subspaces. Although the Hamiltonian of the ring does not commute with the total spin in larger systems, high symmetry may still protect the quantum state and make it more challenging to leak into other subspaces.
III.2 Entanglement entropy for different geometric connectivity
Because tree configuration can lead to a large system size, our current study focuses specifically on the results of the “node” elementary unit’s tree configuration for a system size of $L=8$. In addition to extracting the decoherence lifetime, we also calculate the entanglement entropy, a highly relevant physical quantity in theoretical studies. We partition the system into a target system $\alpha$ and an environment $\beta$, and obtain the density matrix of the target system by performing a partial trace operation: $\rho_{\alpha}(t)={\rm Tr}_{\beta}\rho(t)$, where $\rho(t)=|\Phi(t)\rangle\langle\Phi(t)|$ represents the density matrix of the entire system at time $t$. The entanglement entropy is defined as
$S(t)=-{\rm Tr}_{\alpha}\left[\rho_{\alpha}(t)\ln\rho_{\alpha}(t)\right],$ (4)
where $\rho(t)=|\Phi(t)\rangle\langle\Phi(t)|$ is the density matrix at time $t$.
We investigate the influence of varying the deviation $\sigma$ on the decoherence time for a system with $L=8$ qubits. In the case of the “node” elementary unit, as $\sigma$ increases, the decoherence time gradually decreases for all three connectivity types: ring, chain, and tree. However, the ring geometry shows a faster rate of decrease compared to the chain and tree. Despite the variations, the ring consistently outperforms the chain and tree in terms of decoherence time, indicating that the ring geometry is more stable, in line with our previous conclusions.
Moreover, an important observation from the figures is that the chain and tree geometries exhibit similar decoherence times (Fig. 8). This finding highlights the comparable stability of these two connectivity types. Our results help in strengthening our understanding of how different connectivity patterns impact the overall stability and performance of quantum systems. The geometric configuration plays a crucial role in determining the decoherence behavior, which is valuable for designing robust quantum systems and devising quantum error correction strategies.
Figure 8: The decoherence time $T_{2}^{*}$ as a function of $\sigma$ for elementary units A for ring, chain and tree configuration at system size $L=8$. The interacting strength $J_{0}=100/t_{0}$.
In Fig. 9, we observe a remarkable similarity between the trends of the entanglement entropy and the return probability $P(t)$, indicating that both measures can effectively describe the decoherence time of the system. This finding highlights the potential of utilizing the entanglement entropy as an alternative measure to characterize the degradation of quantum coherence. Additionally, our analysis demonstrates that the stability of the tree configuration configuration is comparable to that of the chain configuration, regardless of whether it is assessed based on the entanglement entropy or the decoherence time. This observation suggests that both chain and tree connectivities exhibit similar levels of vulnerability to decoherence effects, further emphasizing the importance of geometric connectivity in determining the overall stability of quantum systems. These insights contribute to advancing our understanding of the interplay between connectivity patterns and decoherence dynamics, paving the way for improved strategies in quantum information processing and quantum device design.
(a) Refer to caption
(b) Refer to caption (c) Refer to caption
Figure 9: The entanglement entropy $S(t)$ for elementary unit “node” for (a) ring, (b) chain and (c) tree configuration at system size $L=8$. The trend of the entanglement entropy $S(t)$ is very close to that of the return probability $P(t)$, so both can be used to characterize the decoherence lifetime.
IV Conclusions and discussions
In summary, this paper investigates the influence of connectivity on the decoherence time under quasi-static Heisenberg noise. By examining various configurations of elementary units, including rings, chains, and trees, we uncover interesting and counterintuitive results. Specifically, we find that rings exhibit greater stability compared to chains, contrary to the expectation that more links between qubits leads to decreased stability. We attribute this difference to the role of symmetry, which can protect the quantum state from leakage into other subspaces. Additionally, we observe that the “stick” configuration is more stable than the “triangle” configuration, which is expected due to the disparity in their average connectivity. Notably, the disparity in decoherence time between rings and chains is significant, even though their average connectivity is relatively similar. These findings shed light on the complex interplay between connectivity and stability in quantum systems.
Furthermore, our results highlight the comparable stability of tree and chain configurations, as evident from both the entanglement entropy and decoherence time analyses. This insight underscores the importance of considering connectivity patterns when designing quantum systems and devising quantum error correction strategies. Overall, our findings should help in strengthening our understanding of the impact of connectivity on decoherence dynamics, providing inspirations for the development of robust quantum technologies.
Acknowledgements.
This work is supported by the Key-Area Research and Development Program of GuangDong Province (Grant No. 2018B030326001) and the Research Grants Council of Hong Kong (Grant No. CityU 11304920).
Appendix A Dynamic noise
For dynamical noise, we consider the $1/f^{\alpha}$ noise with $\alpha\in[1,3]$, and the time series of the coupling strength $J_{ij}(t)$ is constructed using the fractional Brownian motion (fBm) method Yang and Wang (2016).
(a) Refer to caption (b) Refer to caption
Figure 10: Variation of the decoherence time with the system size $L$ for (a) ring and (b) chain configuration. Each plot represents different values of $\alpha$ ranging from 1 to 3. The decoherence times are rescaled by dividing each value by the maximum $T_{2}^{*}$ for comparison. We take $\sigma=0.5/t_{0}$ and $J_{0}=100/t_{0}$. The elementary unit employed in the study is the “node”. Despite variations in system size and different values of $\alpha$, the decoherence times remain relatively consistent within each geometric connectivity.
We examine the variation of the decoherence time with the system size $L$ for two different geometric connectivities: ring and chain. Fig. 10 shows the results for the ring and chain configurations. Each plot represents different values of $\alpha$, ranging from 1 to 3.
We rescale the decoherence times by dividing each value by the maximum $T_{2}^{*}$ to facilitate comparison. The chosen value of $\sigma$ is 0.5, and the interacting strength is set as $J_{0}=100/t_{0}$. The elementary unit employed in our study is the “node”. It is noteworthy that, despite variations in the system size and different values of $\alpha$, the decoherence times remain relatively consistent within each geometric connectivity. This observation suggests that the chosen geometric connectivity has a more significant impact on the decoherence time than the specific system size or parameter $\alpha$.
Remarkably, irrespective of the system size and different values of $\alpha$, the decoherence times exhibit a consistent pattern within each specific geometric connectivity. This finding underscores the significance of the chosen geometric connectivity in determining the decoherence time dynamics.